### v1

# Is 1:1 Always Most Powerful? Why Unequal Allocation Merits Broader Consideration

Abstract
The principle of allocating an equal number of patients to each arm in a randomized controlled trial is widely accepted as the standard strategy for maximising the trial’s statistical power. However, this long-held belief only holds true if the treatment groups have equal outcome variances, a condition that is often not met in practice. This paper questions the prevalent practice of exclusively defaulting to equal randomisation (ER) and posits that a departure from a 1:1 ratio can be both valid and advantageous. We demonstrate this principle through two simulated case studies, one with a binary endpoint and one with a continuous endpoint, comparing the performance of ER against preplanned Fixed Unequal Randomisation and Response-Adaptive Randomisation targeting Neyman allocation. Our results show that unequal ratios can increase statistical power while simultaneously allocating a substantially larger proportion of patients to the superior treatment arm compared to ER. We conclude that, when unequal variances are suspected, a strategic decision regarding the allocation ratio –- rather than a default 1:1 –- constitutes the superior design choice.
Keywords Wald test, Patient-benefit, Fixed Randomisation, Unequal Randomisation Ratios, Neyman Allocation, Expected Patient Outcomes
1 Introduction
Many trialists and statisticians strongly adhere to the principle that allocating an equal number of patients to each arm in a 1:1 ratio results in the most efficient design, maximising the study’s statistical power. This belief is frequently found in review articles (Dumville et al., 2006; Dibao-Dina et al., 2014; Berger et al., 2021; Freidlin and Korn, 2024), for instance:
“balanced group sizes will maximise a study’s statistical power" (Dumville et al., 2006).
The principle is further reinforced by influential textbooks (Pocock, 2013; Friedman et al., 2015). For example, Chapter 9 of Pocock (2013) details standard power calculations where formulas implicitly assume a 1:1 patient allocation, yielding $n_{C}=n_{T}$ patients per treatment arm (control ($C$) and treatment ($T$)). Building the equal allocation ratio into standard sample size formulae calculations, whether intentional or not, elevates the 1:1 ratio as the default.
The rationale for an equal allocation ratio default is multifaceted (Avins, 1998), reflecting not just statistical considerations but also ethical, financial, and practical ones, including patient recruitment. However, conclusions may be flawed if they rest on the assumption that an equal ratio inherently maximizes power (Torgerson and Campbell, 1997; Wu et al., 2014; Nay et al., 2024). These arguments are often conflated with the ethical and practical justifications being reinforced by the assumption that equal allocation is always the most statistically powerful design:
“Trials using unequal allocation will therefore either have less statistical power or will be more expensive and entail exposing more patients than necessary to a novel intervention and research procedures." (Hey and Kimmelman, 2014).
This long-held assumption that equal allocation maximises power is conditional. Its mathematical validity rests on the crucial prerequisite that the outcome variances within the compared groups are equal. Yet, this assumption of homoscedasticity is often questionable in practice and may often be violated in many settings (Ruxton, 2006). When this condition of equal variances is not met, the optimality of equal allocation is no longer guaranteed. This means that by moving away from a 1:1 ratio, we may achieve superior performance, for instance, through fixed unequal allocation (by sampling more from the more variable arm) or through Response-Adaptive Randomisation (RAR), where allocation probabilities are dynamically adjusted as data accrues and variance estimates improve. The objective of this paper is to challenge the conventional default application of equal allocation and establish a fundamental point: that equal allocation is not a universal optimum. We argue that a well-reasoned allocation ratio, informed by beliefs about outcome variances (even if it deviates from equal allocation), is a superior option to defaulting to 1:1 without careful consideration of outcome group variability.
2 The statistical rationale for increasing power
The allocation ratio that maximises statistical power for a given test statistic depends on:
  1. 1.
    the type of endpoint (binary, discrete, continuous),
  2. 2.
    the measure of interest (e.g., mean difference, odds ratio),
  3. 3.
    and, most critically, the variance of the outcome within each treatment arm.
We illustrate how the above elements interplay when aiming to maximise power using the simple Wald test for comparing two group means. For this test, it has been shown that comparing groups under heteroscedasticity, sampling with an unequal ratio can increase statistical power (Tschuprow, 1923; Neyman, 1934; Robbins, 1952; Rosenberger et al., 2001). By treating the question as a formal optimization problem, one can derive the optimal allocation strategy for the Wald test. This principle, known as Neyman allocation, dictates that to maximize power for a fixed total sample size, the allocation ratio between the treatment and control arms should be proportional to the standard deviations of the outcomes in each group:
$\frac{n_{T}}{n_{C}}=\frac{\sigma_{T}}{\sigma_{C}}$ (1)
Formula (1) demonstrates, that ${n_{T}}={n_{C}}$ is only optimal when variances are equal ($\sigma_{T}=\sigma_{C}$). Otherwise, maximising statistical power necessitates allocating more patients to the arm with higher expected variance.
In the case of binary endpoints, the variance $\sigma_{i}^{2}=p_{i}(1-p_{i})$ within each arm ($i=C,T$) is determined by its success rate, $p_{i}$. Applying the principle of Neyman allocation reveals a critical nuance: ER is only optimal under the null hypothesis ($p_{C}=p_{T}$) or where the success rates are symmetric around 0.5 (i.e., $p_{C}=1-p_{T}$), see Figure 1. This figure also elucidates the possibility of a conflicting situation where maximising statistical power requires the allocation ratio to shift in a direction that favors the currently worse-performing treatment arm.
Figure 1: The horizontal axis shows the varying value of $p_{1}$ while $p_{0}$ takes a fixed value in each sub figure. For each combination of $p_{1}$ and $p_{0}$ the vertical axis shows the value of the theoretical Neyman allocation ($\rho_{N_{1}}$) and Equal Randomisation ($\rho_{\text{ER}}$). Notice that as Neyman allocation is symmetric around 0.5 each sub figure represent two $p_{0}$ values.
This highlights a broader principle: when maximising statistical power, the optimal allocation depends heavily on the context. For instance, with time-to-event data that is exponentially distributed, Neyman allocation always assigns more patients to the superior arm (Baldi Antognini et al., 2020). Furthermore, the choice of effect measure and the statistical test itself alter the optimal ratio (Pin et al., 2024).
More generally, power is optimized by the allocation that maximizes the test statistic’s non-centrality parameter (Hu and Rosenberger, 2003), which for the Wald test, leads to Neyman allocation (Rosenberger et al., 2001). We conclude this section with the theoretical insight that equal allocation can only be expected to be optimal in terms of statistically power only under homoscedasticity; its practical implications will be investigated in the next section.
3 Examples to quantify power improvements
Having established that an equal ratio is not always optimal, we now illustrate just how inefficient it can be. We will do this by exploring the practical implications through two simulated case studies. These simulations compare the statistical power of Equal Randomisation (ER) against a pre-planned Fixed Unequal Randomisation (FUR) and RAR targeting Neyman allocation.
Since our simulated outcomes are independent and identically distributed, our evaluation of ER and FUR is unaffected by the choice of randomisation procedure, so long as the final allocation count is achieved. Our findings therefore apply to any method that achieves the target, including the truncated binomial design, big stick design, permuted block design, and the random allocation rule (Berger et al., 2021). For RAR we use the efficient randomized-adaptive design to target Neyman allocation (Hu et al., 2009).
Case Study 1: early-phase oncology trial
This simulation models a typical early-phase study with a binary endpoint, comparing a new cancer treatment against a placebo in a cohort of 60 patients. For instance, the endpoint could be based on RECIST criteria, where the outcome is dichotomized to 1 (complete/partial response or stable disease) and 0 (disease progression). The expected response rate for the new treatment is 30% ($p_{1}=0.3$), while the placebo response rate is 5% ($p_{0}=0.05$). For our comparisons, the primary analysis uses the Wald test at a two-sided 5% significance level. The FUR design allocates patients to the placebo and treatment groups at a 1:2 ratio. This choice is guided by the optimal Neyman allocation, where $\sigma_{1}/\sigma_{0}\approx 0.678$ (roughly corresponding to a 1:2 ratio). We hypothesize that the observed power gain depicted in Table 1 stems from FUR correcting for the conservativeness of the Wald test under ER in this case.
Table 1: Comparison of allocation strategies for the binary endpoint trial. Metrics include type-I error rate ($p_{0}=p_{1}=0.05$), power, the proportion of patients on the superior arm ($n_{1}/n$), and the Expected Number of Successes (ENS) under the alternative. The Monte-Carlo error for type-I error rate and power is less than $0.2\%$ (Morris et al., 2019).
Design Type-I Error Power $n_{1}/n$ ENS
ER 3.0% 80.0% 0.50 10.5
FUR (1:2) 5.2% 81.5% 0.67 13
As shown in Table 1, moving away from equal allocation increases statistical power and assigns more patients to the superior treatment arm. Using FUR improves power by $1$ percentage point over ER.
Case Study 2: phase III trial with continuous outcomes
The second simulation represents a Phase III trial with a total of 350 patients and a continuous outcome simulated from a normal distribution. A common example of such an endpoint is the change in a symptom score, like systolic blood pressure, from baseline to a follow-up visit. We simulated data from two normal distributions. The control (placebo) arm has a mean of $\mu_{0}=0$, and the treatment arm has a mean of $\mu_{1}=0.5$. The standard deviation in the treatment arm is assumed to be four times higher than in the control arm ($\sigma_{1}=2$, $\sigma_{0}=0.5$). This is consistent with the greater uncertainty and variability typically associated with a new experimental therapy compared to a placebo or an established standard of care.
Table 2: Comparison of allocation strategies for the continuous outcome trial. Metrics include type-I error rate ($\mu_{0}=\mu_{1}=0$), power, the proportion of patients on the superior arm ($n_{1}/n$) and the Expected Mean Response (EMR) under the alternative. The Monte-Carlo error for type-I error rate and power is less than $0.2\%$ (Morris et al., 2019).
Design Type-I Error Power $n_{1}/n$ EMR
ER 5.2% 89.0% 0.50 0.25
FUR (1:2) 4.5% 95.1% 0.67 0.33
FUR (1:3) 4.9% 95.8% 0.75 0.37
FUR (1:4) 5.5% 96.4% 0.80 0.40
In this scenario, we again observe that unequal allocation strategies enhance power and assign a greater proportion of patients to the treatment demonstrating superiority at the trial’s conclusion. As detailed in Table 2, FUR offers a power increase of over 6 to 7 percentage points compared to ER. The closer the patient allocation is to the optimal 1:4 ratio, the greater the statistical power. Consequently, even a partial move from a 1:1 ratio towards 1:4 will yield improvements in power.
4 Discussion
Whether and, crucially, by how much unequal allocation can increase statistical power for a given statistical test depends on several key factors:
  1. 1.
    the endpoint,
  2. 2.
    the total sample size,
  3. 3.
    the magnitude of the treatment effect difference,
  4. 4.
    the relationship between the variances of the treatment arms.
While Neyman allocation might indicate a departure from a 1:1 ratio is optimal, the benefits in terms of statistical power become less pronounced when a study is already robustly powered with ER. This may be the case for studies with a large sample size or an expected large treatment effect. Our key objective, however, was to highlight that implementing unequal allocation is not inherently detrimental to statistical power while it may be highly desirable for other reasons. A reduction in power (compared to ER) would only occur if the chosen allocation ratio diverged from the direction Neyman allocation would prescribe or exceeding it too far, e.g. in Case Study 2 choosing a 2:1 ratio ($76.2\%$ power) or 1:19 ratio ($87.3\%$ power).
Pinpointing the optimal unequal ratio for FUR at the trial’s outset is challenging. This hurdle can be overcome by using RAR instead. The use of RAR to target Neyman allocation in Case Study 1 leads to a more substantial benefit, boosting power by $10$ percentage points. However, using this type of RAR for binary endpoints in combination with the Wald test can lead to serious type-I error inflation. Methods exist to control type-I error when using RAR (Pin et al., 2025). In this case study the RAR design incorporates a “burn-in" period where the first 12 patients are allocated equally, after which the randomization probabilities are adapted. For Case Study 2 with the continuous endpoint and relatively large sample size RAR targeting Neyman does control type-I error and almost reaches the optimal allocation ratio of 1:4 ($n_{1}/n=0.798$ instead of $0.8$). We used a burn-in of 70 patients in total.
While power increased in our examples, FUR or RAR aiming for Neyman allocation can also yield power comparable to ER without significant gains, especially when the optimal allocation is near 1:1. For binary endpoints, Figure 1 shows optimal Neyman allocation deviates little from equal allocation when success probabilities are close to 0.5. Unequal allocation offers the most advantage in terms of statistical power when success probabilities are dissimilar and not close to around 0.5.
It is important to note that our findings are based on a two-arm setting using the parametric Wald test. Choosing an optimal allocation ratio for more specific or complex scenarios and other tests requires further study, (e.g. Jackson and Cox, 2024).
Ultimately, the potential benefit of unequal allocation should be assessed on a case-by-case basis for any given trial design and statistical test. Our analysis demonstrates that considering the Neyman allocation for a study can bring advantages beyond improving statistical power; these gains can translate into significant economic benefits and, most importantly, improved patient outcomes. Consequently, we recommend that trial designs move beyond a default 1:1 ratio. Unequal allocation warrants consideration, as it may offer benefits even when statistical power is on par with a 1:1 approach.
Acknowledgements
  • The members of the working groups listed as affiliations are responsible for the content of this paper. This document does not necessarily represent the consensus or official endorsement of all members of these groups.
  • The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV, DSR, SB), as well as an MRC Biostatistics Unit Core Studentship (LP) and the Cusanuswerk e.V. (LP). SSV is part of PhaseV’s advisory board.
  • The authors acknowledge the use of large language models (LLMs), specifically Gemini (Google), to assist with generating figures and refining grammar and wording in this paper. The LLMs were not used for data analysis, interpretation, or original scientific writing. All content has been carefully reviewed and verified by the authors, who take full responsibility for the integrity and accuracy of the work.